OpenStudy (anonymous):

1. Solve by factoring a. 3x^2-18x=0 b. x(3x-1)=2 c.(4a+1)(3a+1)=13 d.x^2-16x+48=0

OpenStudy (mimi_x3):

a) 3x(x-6) x = 0, 6 b) 3x^2 - x - 2 = 0 (x-1)(3x+2) = 0 x = 1 , -2/3 c) 12a^2 + 7a + 1 = 0 (x+4)(x+3) = 0 x = -4 , -3 d) x^2 -16x+48 = 0 (x-4)(x-12) x = 4 , 12